OpenStudy (anonymous):

Please help me with some chemistry problems? I am so confused. I really need the help.

OpenStudy (anonymous):

1. A cube of an aluminum alloy has a mass of 500.0 grams. Each side of the cube is 5.80 cm long. What is the density of the alloy in g/cm^3? 2. A cook heats a 13.0 kg cast iron skillet (c=0.4521 J/g°C) from room temperature (25.0°C) to 178°C. How many kilojoules of heat are required? 3. If a 200.0 g object needs 2,350 Joules to increase its temperature by 15.0°C, what is its specific heat capacity? 4. The substance decomposed in number 4 above is calcium carbonate, CaCO3. Write the balanced chemical equation for the decomposition. How many moles of oxygen are shown as being produced in the balanced chemical equation? 5. Write the balanced chemical equation for the complete combustion of propane (C3H8). There is one other reactant besides propane. What is it and how many moles of it are required to completely react with one mole of propane? * Thank you SO SO much. I don't want the answers I just want a little help finding it myself. Thank you! Chemistry just doesn't work out for me lol

OpenStudy (aaronq):

post only 1 question per thread. I can help you with the first, 1. A cube of an aluminum alloy has a mass of 500.0 grams. Each side of the cube is 5.80 cm long. What is the density of the alloy in g/cm^3? The formula for density is \(density=\dfrac{mass}{volume}\) so you need the volume (of the cube) and the mass to find the density. For a cube, \(V=side^3\)

OpenStudy (anonymous):

Oh okay thank you

OpenStudy (aaronq):

no problem

OpenStudy (anonymous):

Do you think you might be able to help me with the second one now? I understand if you can't

OpenStudy (aaronq):

sure, but you should post the next question in a new thread. Use: \(q=m*C_p*\Delta T\) q= heat, m=mass, \(C_p\)=specific heat capacity, \(\Delta T\)=change in temp

OpenStudy (aaronq):

just watch your units, you may need to convert